Suppose to copper calorimeter has a mass of 0.1kg and the specific heat capacity of copper C is 400kgk. If the calorimeter contains 0.2kg of water of specific heat capacity 4200kgk and this temperature of the water rises 5oC. What will be the total heat gained?


1
Expert's answer
2020-09-10T11:58:57-0400

According to the equation of the termal balance, the total heat gained is given by the following expression:


"Q = m_cc_c\Delta T + m_wc_w\Delta T"

where "m_c = 0.1kg" is the mass of the calorimeter, "c_c = 400\dfrac{J}{kg\cdot K}" is the specific heat capacity of copper, "m_w = 0.2kg" is the mass of water, "c_w = 4200\dfrac{J}{kg\cdot K}" is the specific heat capacity of water and "\Delta T = 5\degree C = 5K" is the change of the temperature. Thus, obtain:


"Q = (m_cc_c + m_wc_w)\Delta T = (0.1\cdot 400 + 0.2\cdot 4200)\cdot 5 = 4400J"

Answer. 4400 J.